# POD-Attention: Unlocking Full Prefill-Decode Overlap for Faster LLM Inference

### v1

@@ 5.4.2. Scheduling Policy
(a) 2 CTAs per SM.
(b) 4 CTAs per SM.
Figure 13. POD-Attention with varying CTA configs.
@@ 6. Related Work
Optimizing Attention Computation: FlashAttention (flashattention, ) introduced the first specialized implementation of attention, fusing all its operations into a single kernel with tile-based computation. FA-2 (flashattention2, ) improved it further with better work partitioning and load balancing. FlashDecoding (flashdecoding, ) accelerates decode attention by splitting computation along the KV dimension. FlashDecoding++ (flashdecoding++, ) uses asynchronized softmax, double-buffered flat GEMM optimizations, and dataflow-based hardware resource adaptation to accelerate decode. LeanAttention (leanattention, ) follows Stream-K reduction (streamk2024, ) of tiled calculation to enable better load distribution across SMs for decodes. FlashInfer (flashinfer, ) introduced shared-prefix based optimized attention kernels. Compared to works that separately handle prefill and decode, POD-Attention jointly optimizes and fuses them into a single kernel.
FA-3 (flash-attention-3, ) is a recent addition to the FlashAttention family of kernels. It leverages new features available in the NVIDIA Hopper architecture, exploiting the asynchrony of Tensor Cores, the Tensor Memory Accelerator, and the Special Function Units. FA-3 is still under active development at the time of this writing, and many optimizations like ping-pong scheduling were unavailable when POD-Attention was created. We leave extending POD-Attention support to FA-3 and Hopper architecture for future work.
Operation Fusion: Kernel fusion is a commonly used technique for improving GPU performance. Elastic kernels (Elastic_kernel_ASPLOS13, ) proposes restricting resources to enable running multiple kernels concurrently. However, this method provides no guarantee of intra-SM co-location. To overcome this, ISPA (ISPA_TOC23, ) deploys a predetermined number of CTAs for each kernel, less than the number of CTAs that run concurrently on the GPU. Significant a priori profiling is used to determine the appropriate CTA sizes to allow for both kernels to execute concurrently. This can be tedious for attention kernels with dynamically changing input sizes, and makes load balancing between the prefill and decode operations difficult, as one operation completing early leaves resources underutilized. HFuse (HFuse_CGO_2022, ) fuses operations in warp-parallel fashion, providing source-to-source compilation tools to fuse kernels. SM-centric scheduling (smcentric:ICS:2015, ) uses the SM counter to assign work to CTAs, which we leverage in POD-Attention. We do not consider NVIDIA support for MPS (nvidia:mps, ) and MIG (nvidia:mig, ) in this paper since these features cater to multi-process scenarios which is different from our use case.
@@
Optimizing LLM Inference: Optimizing LLM serving systems is an active area of research  (sarathiserve2024, ; splitfuse2024, ; patel2023splitwise, ; sheng2023fairness, ; moedeployment, ; vllmsosp, ; orca, ; dynamollm2024, ; fastserve, ; loongserve2024, ; powerinfer2023, ). Orca  (orca, ) introduced iteration-level scheduling to eliminate compute fragmentation when requests of different lengths are batched together. PagedAttention (vllmsosp, ) and vAttention (vattention2024, ) proposed different techniques for dynamic memory management for LLM inference. Sarathi-Serve (sarathiserve2024, ) leverages chunked prefills to enable stall-free batching. In contrast, Splitwise (patel2023splitwise, ), DistServe (distserve2024, ) and TetriInfer (tetriinfer, ) disaggregate the prefill and decode phases onto different GPU nodes to avoid interference between these phases. Various recent works have also proposed overlapping compute with communication to improve resource utilization (coconetasplos, ; flux2024, ; googleoverlap, ).
@@
Similar to POD-Attention, NanoFlow (nanoflow2024, ) also targets improving intra-device resource utilization, albeit with a contrasting approach. NanoFlow divides a batch into smaller operation-level nano-batches and schedules them in a way that overlaps operations with complementary resource profiles via CUDA streams. In contrast, POD-Attention tries to maximize resource-utilization within a given batch by fusing prefill and decode attention computation. While NanoFlow requires large batch sizes in order to benefit from batch splitting, POD-Attention is useful when attention consumes a significant amount of time. Therefore, NanoFlow seems more suitable for small-context scenarios whereas POD-Attention targets long-context scenarios that depend on hybrid batching for efficient LLM serving.
@@ 7. Conclusion
We introduce POD-Attention — the first attention kernel specialized to compute prefill and decode attention in parallel such that both compute and memory bandwidth can be utilized simultaneously. POD-Attention enables efficient hybrid batching based LLM inference by accelerating attention computation by up to $75\%$ (mean 28%) compared to using independently optimized prefill and decode attention kernels. POD-Attention also improves the end-to-end serving throughput by up to $22\%$, while significantly reducing latency over state-of-the-art LLM serving systems Sarathi-Serve and vLLM.

### v2

@@ 5.4. Sensitivity Studies
(a) 2 CTAs per SM.
(b) 4 CTAs per SM.
Figure 13. POD-Attention with varying CTA configs.
@@ 6. Related Work
Optimizing GPU execution and LLM serving systems is an active area of research (Yu et al., 2022; Kwon et al., 2023; Agrawal et al., 2024c; Prabhu et al., 2025; Sanovar et al., 2024; Zhu et al., 2024; Zhong et al., 2024; Hu et al., 2024; Holmes et al., 2024; Patel et al., 2024; Stojkovic et al., 2024; Wu et al., 2024; Agrawal et al., 2024a; Huang et al., 2024; Cao et al., 2024; Kang et al., 2024; Zheng et al., 2024; Jangda et al., 2024).
Optimizing Attention Computation: FlashAttention (Dao et al., 2022) introduced the first specialized implementation of attention, fusing all its operations into a single kernel with tile-based computation. FA-2 (Dao, 2024) improved it further with better work partitioning and load balancing. FlashDecoding (Dao et al., 2023) accelerates decode attention by splitting computation along the KV dimension. FlashDecoding++ (Hong et al., 2024) uses asynchronized softmax, double-buffered flat GEMM optimizations, and dataflow-based hardware resource adaptation to accelerate decode. LeanAttention (Sanovar et al., 2024) follows Stream-K reduction (Osama et al., 2023) of tiled calculation to enable better load distribution across SMs for decodes. FlashInfer (Ye et al., 2025) introduced shared-prefix based optimized attention kernels. Compared to works that separately handle prefill and decode, POD-Attention jointly optimizes and fuses them into a single kernel.
FA-3 (Shah et al., 2024) is a recent addition to the FlashAttention family of kernels. It leverages new features available in the NVIDIA Hopper architecture, exploiting the asynchrony of Tensor Cores, the Tensor Memory Accelerator, and the Special Function Units. FA-3 was under active development at the time of writing this paper and hence we leave extending POD-Attention support to FA-3 and Hopper architecture for future work.
Operation Fusion: Kernel fusion is a commonly used technique for improving GPU performance. Elastic kernels (Pai et al., 2013) proposes restricting resources to enable running multiple kernels concurrently. However, this method provides no guarantee of intra-SM co-location. To overcome this, ISPA (Zhao et al., 2023) deploys a predetermined number of CTAs for each kernel, less than the number of CTAs that run concurrently on the GPU. Significant a priori profiling is used to determine the appropriate CTA sizes to allow for both kernels to execute concurrently. This can be tedious for attention kernels with dynamically changing input sizes, and makes load balancing between the prefill and decode operations difficult, as one operation completing early leaves resources underutilized. HFuse (Li et al., 2022) fuses operations in warp-parallel fashion, providing source-to-source compilation tools to fuse kernels. SM-centric scheduling (Wu et al., 2015) uses the SM counter to assign work to CTAs, which we leverage in POD-Attention.
@@
Optimizing LLM Inference: Optimizing LLM serving systems is an active area of research  (Agrawal et al., 2024c; Holmes et al., 2024; Patel et al., 2024; Sheng et al., 2024; Huang et al., 2024; Kwon et al., 2023; Yu et al., 2022; Stojkovic et al., 2024; Wu et al., 2023, 2024; Song et al., 2024). Orca  (Yu et al., 2022) introduced iteration-level scheduling to eliminate compute fragmentation when requests of different lengths are batched together. PagedAttention (Kwon et al., 2023) and vAttention (Prabhu et al., 2025) proposed different techniques for dynamic memory management for LLM inference. Sarathi-Serve (Agrawal et al., 2024c) leverages chunked prefills to enable stall-free batching. In contrast, Splitwise (Patel et al., 2024), DistServe (Zhong et al., 2024) and TetriInfer (Hu et al., 2024) disaggregate the prefill and decode phases onto different GPU nodes to avoid interference between these phases. Various recent works have also proposed overlapping compute with communication to improve resource utilization (Jangda et al., 2022; Chang et al., 2024; Wang et al., 2022).
@@
Similar to POD-Attention, NanoFlow (Zhu et al., 2024) also targets improving intra-device resource utilization, albeit with a contrasting approach. NanoFlow divides a batch into smaller operation-level nano-batches and schedules them in a way that overlaps operations with complementary resource profiles via CUDA streams. In contrast, POD-Attention tries to maximize resource-utilization within a given batch by fusing prefill and decode attention computation. While NanoFlow requires large batch sizes in order to benefit from batch splitting, POD-Attention is useful when attention consumes a significant amount of time. Therefore, NanoFlow seems more suitable for small-context scenarios whereas POD-Attention targets long-context scenarios that depend on hybrid batching for efficient LLM serving.
@@ 7. Conclusion
We introduce POD-Attention — the first attention kernel specialized to compute prefill and decode attention in parallel such that both compute and memory bandwidth of a GPU can be utilized simultaneously. POD-Attention enables efficient hybrid batching based LLM inference by accelerating attention computation by up to $59\%$ (mean 28%) compared to using independently optimized prefill and decode attention kernels. POD-Attention also improves the end-to-end serving throughput by up to $22\%$, while significantly reducing latency over state-of-the-art LLM serving systems Sarathi-Serve and vLLM.
Acknowledgements.
We thank our shepherd Tim Rogers, the anonymous ASPLOS reviewers and Ajay Nayak for their valuable feedback on various aspects of the paper. We also thank Zihao Ye for various helpful discussions on POD-Attention and FlashInfer. Aditya K Kamath and Simon Peter are supported by National Science Foundation grant CNS-2212580.
Appendix A Artifact Appendix
A.1. Abstract
POD-Attention is a GPU kernel that overlaps prefill and decode attention operations for large language models. POD-Attention is built on top of FlashAttention kernels (v2.6.1) (Dao, 2024) and is integrated with Sarathi-Serve (Agrawal et al., 2024c) – a state-of-the-art hybrid batching based LLM inference scheduler.
A.2. Artifact check-list (meta-information)
  • Compilation: CUDA 12.4, GCC 11.4.
  • Model: Llama-2-7B (lla, 2024a), Llama-3-8B (lla, 2024b), Yi-6B (yi-, 2024).
  • Data set: arXiv-Summarization (arx, 2024).
  • Run-time environment: Ubuntu 22.04, CUDA 12.4, Python 3.12, and PyTorch 2.4.
  • Hardware: 1–2 NVIDIA A100 80 GB GPUs, x86 machine.
  • How much time is needed to prepare workflow?: 1 minute with Docker image. 1–2 hours if installing from source.
  • How much time is needed to complete experiments (approximately)?: Approx. 18 hours.
  • Publicly available?: Yes.
  • Archived (provide DOI)?: 10.5281/zenodo.14770841
A.3. Description
A.3.1. How to access
We provide the source code in various forms: Docker container (see A.3.3), GitHub repository (https://github.com/microsoft/vattention/tree/main/pod_attn), and Zenodo (https://doi.org/10.5281/zenodo.14770840).
A.3.2. Hardware dependencies
This artifact requires an x86 machine with 2 NVIDIA A100 GPUs with 80GB memory each. If only one GPU is available, all experiments can be conducted in full, except for Table 6 and the results for Llama-2-7B and Llama-3-8B in Figure 12.
A.3.3. Software dependencies
POD-Attention has been tested on a machine with Ubuntu 22.04. All other software dependencies are resolved while installing.
A.3.4. Data sets
Some experiments are based on the arXiv-Summarization dataset. We use a subset of the dataset available in the traces/ folder of the artifact.
A.3.5. Models
This artifact evaluates Yi-6B, Llama-2-7B and Llama-3-8B. Accessing Yi-6B and Llama-2-7B is straightforward but accessing Llama-3-8B requires logging into huggingface with the user’s private token (HF_TOKEN below): {Verbatim}[frame=single,rulecolor=,fontfamily=zi4,fontsize=] $huggingface-clilogin--tokenHF_{T}OKEN\par$
A.4. Installation
We provide two methods of installing and testing: using Docker (recommended) or manual installation.
A.4.1. Docker installation (recommended)
We provide a docker image for POD-Attention with all its dependencies pre-installed. You can launch the docker container and navigate to the artifact directory as follows: {Verbatim}[frame=single,rulecolor=,fontfamily=zi4,fontsize=] $dockerrun--gpusall-it\ -p8181:8181--rm--ipc=host--cap-add=SYS_{A}DMIN\ rnp1910/pod_{a}ttention:asplos_{2}5_{p}ytorch_{r}un$ cd /workspace/vattention/pod_attn
A.4.2. Manual installation
For manual installation, we can download POD-Attention (available in vAttention repository) to home directory to install it. We use Anaconda for the appropriate versions of CUDA, Python, and PyTorch. This can take up to 2 hours.
\$ git clone \
https://github.com/microsoft/vattention.git
\$ cd vattention/pod_attn/
# Install miniconda; skip if already installed
\$ make install_miniconda
\$ bash # Refresh shell and activate
\$ conda activate pod_attn
# Install CUDA Toolkit
(pod_attn)\$ conda install -y -c \
conda-forge cuda-toolkit=12.4.0
# Install dependencies
(pod_attn)\$ pip install -r requirements.txt
(pod_attn)\$ pip install flashinfer==0.1.5 \
-i https://flashinfer.ai/whl/cu124/torch2.4
# Install POD-Attention and vAttention
(pod_attn)\$ make install_all
A.5. Experiment workflow
The source code for POD-Attention kernel is available in the vattention/pod_attn/ folder. Our evaluation primarily contains two kinds of experiments: attention performance (Figures 1, 6, 10, 11, 13, 14) and end-to-end LLM performance (Figure 12 and Table 6). Figure 7 evaluates various kernel fusion strategies with a micro-benchmark. Most of these require only one GPU except for Table 6 and Figure 12 (for Llama-2-7B and Llama-3-8B) that require two GPUs. Use the Makefile present in the vattention/pod_attn/ folder to run experiments as follows:
make figure1 # 2 minutes; sudo used by script
make figure6 # 2 minutes
make figure7 # 2 minutes
make figure10 # 1 minute; sudo used by script
make figure11 # 2 hours
make figure12 # 9 hours
make figure13 # 1 minute
make figure14 # 1 minute
make table6 # 4 hours
A.6. Evaluation and expected results
The artifact scripts redirect the raw output numbers and logs to output/ folder, while the plotted graphs can be found in the graphs/ folder. Tables are saved as CSVs in the same folder. Results may have minor runtime variations from those reported in in the paper, but general trends should hold.